If X is a R.V with PDF f(x) = x/6 when x = 1, 2, 3 and f(x) = 0, otherwise. Find E(X+2)2

1
Expert's answer
2021-09-29T00:29:41-0400
"\begin{matrix}
 (x+2)^2 & 9 & 16 & 25 \\
 p & 1/6 & 2/6 & 3/6
\end{matrix}"

"E[(X+2)^2]=9(\dfrac{1}{6})+16(\dfrac{2}{6})+25(\dfrac{3}{6})=\dfrac{58}{3}"

"E[(X+2)^2]=\dfrac{58}{3}"